OpenStudy (anonymous):

the age of a father is twice the square of his sons age 8 yr hence the age of father willl be 4 yrs more then three times the age of his son calculate their present age

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

im gettigf 2x^2 - 3x - 20 = 0

OpenStudy (anonymous):

@hartnn

OpenStudy (godlovesme):

Want r u supposed to solve for?

OpenStudy (anonymous):

thier present age

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Hmm.. Confusing. @ganeshie8 @mathmate @SolomonZelman @jhonyy9 @paki Any ideas?

OpenStudy (anonymous):

2X^2 - 3X - 20 = 0 AB

OpenStudy (anonymous):

Let f= father's present age Let s= son's present age \[f=2(s)^{2}\] Eight year hence \[f+8=2(s)^{2}+8\] So we now have this equation \[2(s)^{2}+8=3(s+8)+4\]\[2s ^{2}+8=3s+24+4\]\[2s ^{2}+8=3s+28\]\[2s ^{2}-3s-20=0\] Solve for s using the quadratic formula \[s=4\] So son is 4 years old and father is 32 years old.